OpenStudy (anonymous):

how many different arrangements can be made with the letters in the word SCARLET

OpenStudy (anonymous):

7 factorial

OpenStudy (lgbasallote):

SCARLET has 7 letters. no letter repeats so it's just 7!

OpenStudy (anonymous):

These are the options

OpenStudy (lgbasallote):

\[7! = 7 \times 6 \times 5 \times 4 \times 3 \times 2\] use a calculator

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

Its k

OpenStudy (lgbasallote):

welcome welcome

OpenStudy (anonymous):

I am new here so how do i ask more questions?